OpenStudy (anonymous):

please answer with solution find the radius of a sphere whose surface area and volume are numerically equal

OpenStudy (lgbasallote):

formula for surface area of a sphere is \[SA = 4\pi r^2\] the formula for volume is \[\frac 43 \pi r^3\] if they are equal then \[4\pi r^2 = \frac 43 \pi r^3\] if i divide both sides by \(4\pi\) i get \[\implies r^2 = \frac 13 r^3\] i multiply both sides by 3 \[\implies 3r^2 = r^3\] i subtract both sides by 3r^2 \[\implies r^3 - 3r^2 = 0\] factor out r^2 \[\implies r^2 ( r - 3) = 0\] solve for r now does that help?

OpenStudy (unklerhaukus):